A portion of the stockholders’ equity section from the balance sheet of Walland Corporation appears as follows.

Stockholders' equity:
Preferred stock, 9% cumulative, $50 par, 31,000 shares authorized, issued, and
outstanding
$ 1,550,000
Preferred stock, 12% noncumulative, $100 par, 7,200 shares authorized, issued, and
outstanding
720,000
Common stock, $5 par, 440,000 shares authorized, issued, and outstanding 2,200,000
Total paid-in capital $ 4,470,000

Assume that all the stock was issued on January 1, Year 1 and that no dividends were paid during the first two years of operation. During the third year, Walland Corporation paid total cash dividends of $736,000.

a. Compute the amount of cash dividends paid during the third year to each of the three classes of stock.

b. Compute the dividends paid per share during the third year for each of the three classes of stock. (Round your answers to 2 decimal places.)

c. What was the average issue price of each type of preferred stock?

a. Cumulative preferred stock
Noncumulative preferred stock
Common stock
b. Cumulative preferred stock per share
Noncumulative preferred stock per share
Common stock per share
c. Cumulative preferred stock per share
Noncumulative preferred stock per share
Answer #1

a. Compute the amount of cash dividends paid during the third year to each of the three classes of stock.

Preferred cumulative = 31,000*50*9%*3 = 418,500

Preferred non cumulative = 7,200*100*12% = 86,400

Common stock = 736,000-418,500-86,400 = 231,100

b. Compute the dividends paid per share during the third year for each of the three classes of stock.

Preferred cumulative = 418,500/31,000 = 13.5

Preferred non cumulative = 86,400/7200 = 12

Common stock = 231,100/440,000 = 0.53

c. What was the average issue price of each type of preferred stock?

Preferred cumulative = 1,550,000/31,000 = 50

Preferred cumulative = 720,000/7200 = 100